OpenStudy (vampirediaries):

What is the inequality for the graph? I just need someone to check to see if my answer is right.

OpenStudy (vampirediaries):

\[y \ge -x+6\] or \[y \ge-\left| x-5 \right|+1\]

OpenStudy (igreen):

It's obviously not the first option..otherwise the outline would be a straight line.

OpenStudy (igreen):

@vampirediaries

OpenStudy (vampirediaries):

True. What about the second inequality?

OpenStudy (vampirediaries):

@iGreen

OpenStudy (igreen):

You only have 2 choices..don't you?

OpenStudy (igreen):

Well the second one looks incorrect also..

OpenStudy (vampirediaries):

It's an open-ended question

OpenStudy (vampirediaries):

@iGreen

OpenStudy (igreen):

What do you mean..?

OpenStudy (vampirediaries):

I don't have any answer choices to pick from, so I have to create an equality

OpenStudy (igreen):

Oh, okay.

OpenStudy (vampirediaries):

Can you help me figure out the equality?

OpenStudy (igreen):

Is the line dotted or solid..?

OpenStudy (igreen):

It looks solid to me.

OpenStudy (vampirediaries):

Yeah, it's a solid line

OpenStudy (igreen):

Okay, the inequality is going to be: \(y \ge |x - 5| + 1\)

OpenStudy (igreen):

You can graph it here to see for yourself: https://www.desmos.com/calculator

OpenStudy (igreen):

Just copy and paste it into a box on the left.

OpenStudy (igreen):

I just experimented with the graph..

OpenStudy (igreen):

Well subtracting the x in the brackets shifts it to the right, and adding it shifts it to the left. Adding outside of the bracket makes it go up, and subtracting outside of the bracket makes it go down.

OpenStudy (igreen):

Therefore, you subtract x in the bracket by 5, and you add outside of the bracket by 1.

OpenStudy (vampirediaries):

Ok. So, what about the y-intercept? y=6

OpenStudy (igreen):

Yeah, the y-intercept is 6.

OpenStudy (igreen):

@vampirediaries

OpenStudy (vampirediaries):

So I just leave it out of the inequality?